# Equal Contribution

### v2

@@ Abstract
1 Introduction
Creating 3D objects that facilitate human interactions is a critical and long-standing problem in industrial engineering, with applications ranging from manufacturing to developing simulation environments. With the recent proliferation of AI-facilitated 3D content creation [51, 34, 11, 71, 84, 39, 54, 38, 40], a line of efforts has been made to model interactable objects, particularly articulated ones, with 3D generative models [29, 37]. Imagine having a cabinet with a broken drawer; rather than engaging in labor-intensive carpentry with jigs and saws, a more efficient solution would be to generate a 3D model of the drawer and print it using a 3D printer, thus eliminating the need for specialized human expertise.
However, unlike rigid objects, whose modeling primarily focuses on shapes and appearances, interactable object modeling presents the additional challenge of intricate part motions under human interactions, constrained by geometry and physics. This complexity is particularly problematic for real-world applications, where high accuracy is essential. For instance, in the case of a drawer, even a slight size mismatch or a minor surface bump can cause it to get stuck in the cabinet.
In this paper, we address the problem of generating 3D object parts that are physically plausible both in their rest states and under interactions. Previous work on physics-constrained shape generation has predominantly concentrated on rigid objects, emphasizing connectivity [43, 22, 23] or structural stability [42]. However, a challenge arises for interactable objects with part motions: while the generated shapes are static, the inherent motion state is also crucial for physical plausibility. Thus, constraints on both rest and motion states must be carefully considered.
For evaluation, we construct a benchmark with metrics for assessing both the shapes and the physics of the generation results. Besides the widely adopted metrics for 3D generative models, we introduce a physical plausibility metric that directly evaluates the success rates of part motions. With both qualitative and quantitative results, we show that our framework generates shapes of high quality, outperforming methods that lack proper modeling of geometric conditioning or physical constraints. Furthermore, we demonstrate two downstream applications: 3D printing and robot manipulation, showcasing our strong potential in real-world applications that require high physical plausibility.
To summarize, our key contributions are:
  • We introduce the task of physically plausible part completion for interactable objects and provide a benchmark with holistic evaluation metrics for both shape and physics.
  • We propose a novel diffusion-based part generative model; it incorporates geometric conditioning and physical constraints by formulating them as stability and mobility losses, respectively.
  • Experimentally, we show compelling results in both shape quality and physical plausibility compared with baseline models, demonstrating the efficacy of our model.
  • We showcase physics-demanding applications of our model, including 3D printing, robot manipulation, and sequential part generation, where high physical plausibility is crucial to the task.
2 Related Work
Articulated Object Modeling. Articulated object modeling is a crucial and longstanding field in 3D vision and robotics, encompassing a wide range of work in perception [17, 74, 12, 31, 35, 16, 18, 15, 19], reconstruction [2, 46, 25, 67], and generation [29, 37]. These contributions have greatly advanced applications in simulation [19, 73] and robot manipulation[4, 13, 68, 75, 32, 72, 30, 15, 16]. Beyond these perspectives, the intricate per-part state changes and inter-part motions, which formulate how objects interact and respond to interactions, also play important roles. For instance, [17] categorizes parts based on their actions, [74] learns part segmentation by analyzing point clouds during part motions, and [67, 46, 25] focus on reconstructing parts and joints across various motion states. However, most studies primarily focus on part motions at the pose level, often neglecting the critical interplay of underlying physics with detailed part geometry. This oversight is particularly significant in 3D generation, where physical plausibility is essential for the practical application of generated models. Nonetheless, this critical aspect of modeling remains under-explored in recent studies, which focus more on structure and shape [29, 37]. We direct readers to survey [36] for a more comprehensive overview.
Physics-Aware Shape Generation. Previous research in physics-aware 3D object generation has primarily focused on rigid shapes without complex self-interactions [43, 22, 23, 42]. For instance, [43, 22, 23] have used persistence diagrams to ensure topological connectivity, and [42] introduced a differentiable physical simulation layer to enhance stability under gravity. In contrast, physical principles have been extensively applied in other areas, such as in PhyScene [73], which aids in synthesizing interactable 3D scenes, and in human pose and motion generation where studies [80, 81, 69] incorporate physics-aware constraints to reduce collisions and enhance realism. Additionally, built on diffusion models, SceneDiffuser [24] and PhysDiff [77] achieve physically plausible generation by integrating physical optimization guidance into the denoising process. Our work uniquely extends these principles by embedding physical constraints, including stability and mobility, into a generative diffusion model, enabling the generation of physically plausible parts for interactable articulated objects.
3D Shape Completion. Shape completion is a vital step for reconstructing the missing parts of 3D shapes, traditionally tackled with techniques such as Laplacian hole filling [61, 48, 82] and Poisson surface reconstruction [26, 27]. These methods primarily address small gaps and basic geometric forms. Another strategy employs structural regularities, like symmetries, to infer unobserved parts of shapes [66, 44, 50, 55, 62]. With the advent of extensive 3D datasets, retrieval-based approaches [64, 33, 47, 28] have emerged, searching databases to find the best matches for incomplete inputs, alongside learning-based methods [49, 14, 10, 8, 76, 20, 60, 6] that minimize discrepancies between network predictions and actual shapes. Notably, 3D-EPN [10] utilizes a 3D encoder-decoder architecture to predict complete shapes from partial volumetric data, while Scan2Mesh [9] transforms range scans into 3D meshes through direct mesh surface optimization. PatchComplete [53] leverages local structural priors to fill in shapes from unseen categories. Alternatively, generative methods like GANs [83, 3, 70, 79, 56] and AutoEncoders [45, 1] offer a different approach by generating diverse, plausible shapes from partial inputs, though they often compromise on completion accuracy. In contrast, our method leverages diffusion models and introduces physics-aware loss guidance, effectively reducing surface artifacts and surpassing current state-of-the-art methods in producing realistic and physics-plausible 3D shapes, allowing applications like 3D printing and robot manipulation.
3 Method
Given the full point cloud of an articulated object $\mathcal{O}$ with a missing part of a known category (e.g., drawer), we aim to generate a replacement part $p$ that integrates seamlessly into the object while adhering to geometric and physical constraints. We employ volumetric Signed Distance Fields (SDF) to model object parts and train a 3D-VQVAE to compress these models into a discrete and compact latent space. Unlike prior research that employs SDF representation overlooks the object’s scale as generating SDF usually involves normalizing the part to the unit cube, our framework also emphasizes accurately scaling the objects for manufacturing purposes. Upon this part representation, our method utilizes a coarse-to-fine approach, involving two diffusion-based sub-models: a pose proposal model that determines the part’s approximate position at the coarse level, and a part code generator that refines the shape based on geometric conditions (Sec. 3.1) and physical constraints (Sec. 3.2). This integrated framework ensures that the generated part can interact properly with external forces and fit accurately within the existing object structure. An overview of our pipeline is shown in Fig. 2.
Part-Pose Proposal. We propose a diffusion-based pose estimator to locate the bounding box of the part to be generated given the object point cloud $\mathcal{P}$. Inspired by GenPose [78], we train a score-based diffusion model $\Phi_{\theta}$ and an energy-based diffusion model $\Psi_{\phi}$. Then we first generate pose candidates $\{p^{i}\}_{i=1}^{K}$ from $\Phi_{\theta}$ and then compute the pose energies $\Psi_{\phi}(p^{i},\mathcal{P})$ for candidates via $\Psi_{\phi}$. After that, by ranking the candidates with the energies and then filtering out low-ranking candidates, the remaining bounding box candidates are aggregated into the final output using mean pooling.
3.1 Geometry-Conditioned Part Generation
@@
SDF Generation with Latent Diffusion. SDFusion [5] introduces a diffusion-based 3D generative model. The approach begins by compressing high-dimensional SDFs into a compact latent space using a 3D-VQVAE. Specifically, an SDF $X\in\mathbb{R}^{D\times D\times D}$ is encoded into a latent vector $z\in\mathbb{R}^{d\times d\times d}$ by the encoder $E_{\phi}$. The encoded SDF is then reconstructed by the decoder $D_{\tau}$ after passing $z$ through a vector quantization codebook VQ, i.e.,
@@
Afterward, a conditional diffusion model is trained on the learned latent distribution. Techniques including the classifier-free guidance [21] and DDIM [58] sampling are utilized to enhance the flexibility and efficiency of the generation process.
Geometric Conditioning. The above framework allows the generation of missing parts using point cloud as the condition. However, due to limited training data and the point cloud’s inherent ambiguity, relying solely on the point cloud often yields unsatisfactory results. Thus, we leverage the part bounding box predicted by our pose estimation model as an additional condition. Specifically, given the predicted bounding box of the desired part, we use the ratio of the dimensions along the three axes as an auxiliary condition for the diffusion model. The length ratio rather than absolute values is used since SDF representation is scale-agnostic. Note that although we allow rotation to the input point cloud, the mesh is always generated in its canonical pose.
Empirically, we adopt the PointNet++ [52] ($\phi_{1}$) to encode the point cloud $c_{1}\in\mathbb{R}^{N\times 3}$, and a sinusoidal positional encoding followed by a two-layer MLP ($\phi_{2}$) to encode the bounding box ratio $c_{2}\in\mathbb{R}^{3}$. Then, we concatenate two control signals then inject them into the model using self-attention layers. The training objective for the diffusion model $\theta$ and encoders $\phi_{1},\phi_{2}$ is formulated as:
$L(\theta,\phi_{1},\phi_{2}):=\mathbb{E}_{z,c,e,t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t},t,D(\phi_{1}(c_{1})),D(\phi_{2}(c_{2})))\right\|^{2}\right],$ (2)
where $D(\cdot)$ denotes a dropout operation, $c_{1},c_{2}$ are the conditions, and $z_{t}$ is the noisy latent variable at timestamp $t$. Please refer to Appendix 0.A for details on model architecture.
@@
3.2 Loss-Guided Physical Constraints
Given an object’s point cloud, we predict the bounding box for the missing part, generate the desired part by sampling from the diffusion model using the object’s point cloud and the bounding box as conditions, and scale the generated mesh using the bounding box volume ratio. However, this generation process is agnostic to physical constraints and may result in parts incompatible with the original object. To address this, we introduce physics-aware losses to guide the diffusion sampling process, ensuring physical compatibility.
Stability Losses. The most critical criterion for the generated part is to avoid collision with the provided object point cloud $\mathcal{P}$. Therefore, we define the collision loss by considering the predicted part-SDF value $X(p)$ of the object point $p$ as follows:
$\ell_{collision}(X,\mathcal{P})=\sum_{p\in\mathcal{P}}\text{ReLU}(-X(p)-\alpha),$ (3)
@@ 3.2 Loss-Guided Physical Constraints
$\ell_{contact}(X,\mathcal{P})=\min_{p\in\mathcal{P}}\text{ReLU}(X(p)-\alpha),$ (4)
@@
$\ell_{collision-m}(X,\mathcal{P})$ $=\mathbb{E}_{\hat{X}\sim M(X)}\sum_{p\in\mathcal{P}}\text{ReLU}(-\hat{X}(p)-\alpha),$ (5)
$\ell_{contact-m}(X,\mathcal{P})$ $=\mathbb{E}_{\hat{X}\sim M(X)}\min_{p\in\mathcal{P}}\text{ReLU}(\hat{X}(p)-\alpha).$ (6)
@@
Loss-Guided Sampling. Conditional diffusion models aim to generate samples $x_{0}$ conditioned on a specific criterion $y$, where the conditional score at timestamp $t$ is derived using Bayes’ rule:
@@
$\overline{\epsilon}_{\theta}(x_{t},t,y)=\epsilon_{\theta}(x_{t},t)-\sqrt{1-\overline{\alpha}_{t}}\,\omega\nabla_{x_{t}}\log{p_{t}}(y|x_{t}),$ (8)
where the gradient term $\nabla_{x_{t}}\log{p_{t}}(y|x_{t})$ is written as $-\frac{1}{\sqrt{1-\overline{\alpha}_{t}}}(\epsilon_{\theta}(x_{t},t,y)-\epsilon_{\theta}(x_{t},t))$ in classifier-free guidance [21].
To introduce physical constraints into the sampling process, we approximate the gradient using our defined physics-aware loss functions [7, 59]. As loss functions are defined on noise-free samples, we compute the loss on the predicted noise-free sample at timestamp $t$, denoted as $\hat{x}_{0}$:
$\nabla_{x_{t}}\log{p_{t}}(y|x_{t})$ $\approx\nabla_{x_{t}}\log{p_{t}}(y|\hat{x}_{0})$ (Eq. (15) in [7])
$=\nabla_{x_{t}}\log\frac{\exp(-\ell_{y}(\hat{x}_{0}))}{Z}=-\ell_{y}(\hat{x}_{0}).$ (9)
Finally, we combine classifier-free guidance and loss guidance to obtain the modified score:
@@
$\ell_{y}(x_{0})=\ell_{collision}+\ell_{contact}+\ell_{collision-m}+\ell_{contact-m}.$ (11)
4 Experiments
This section demonstrates the effectiveness of our method in generating four types of common articulated objects: slider drawers, hinge doors, hinge knobs, and line handles, where the first two allow self-movements and the latter two must be attached to other parts (denoted as dependent parts). We then explore its potential applications in 3D printing, robot manipulation, and sequential part generation tasks. Appendix 0.B presents more experimental results.
Data and Implementation. We train our models on the GAPartNet dataset [17], a part-centric object model dataset with part-level annotations and actionability alignment across part classes. Approximately 90% of instances are allocated for training for each part category, with the remainder reserved for evaluation. All part categories utilize the same VQVAE model, trained across all part categories at an input SDF resolution of 128, while a separate diffusion model is tailored for each category. To ensure comparability, both our method and all baselines involve random rotation of the input point clouds around the vertical axis during training. For testing, we sample three different rotation angles for each object instance. Further implementation details are available in Appendix 0.A.
Baselines. Our primary baseline is the Conditional 3D Variational Autoencoder (3D-CVAE) [57], which learns a VAE on part SDF with the object’s point cloud as the condition. To ensure fair comparisons, we scale the generated parts using our predicted pose and then use the same settings to process the results. To further demonstrate the effectiveness of our loss guidance, we also compare the vanilla SDF diffusion without our proposed loss functions.
4.1 Physics-Plausible Generation
Metrics. We introduce a physical plausibility metric to evaluate the viability of generated shapes. Specifically, we first pre-process the meshes, excluding those with multiple large connected parts or those not watertight as physically infeasible. For the remaining meshes, a grid search around the part’s ground truth position is conducted to locate positions where both collision loss (Eq. 5) and contact loss (Eq. 6) fall below a specified threshold. A part is deemed physically plausible if its movement along the $+x$ and $-x$ axes (for slider drawers) or along the $-z$ axis (for other part categories) causes these losses to exceed the threshold. Notably, for hinge doors, a shared rotation axis is used during loss-guided sampling, while the ground truth rotation axis is used for evaluation. Please see Sec. 0.A.1 for a comprehensive explanation and pseudo-code. Besides the physical plausibility metric, we also utilize Chamfer Distance (CD) and F-score (@2%) [65] for evaluation.
Self-moving Part Generation. As shown in Tab. 1, our method significantly outperforms the baseline in terms of generation quality (CD and F-score) and physical plausibility in generating self-moving parts. The results also highlight the essential role of our proposed physical guidance in the generation process. Notably, as illustrated in Fig. 4, while the baseline method or the absence of physical guidance often leads to collisions or misalignments at the interface between the object and the generated part, our approach consistently delivers parts that seamlessly integrate with the original object, establishing a robust foundation for subsequent dependent part generation.
Dependent Part Generation. As shown in Tab. 2, our approach remains superior to baseline methods dependent parts. Fig. 5 reveals that while the baseline method frequently produces a degenerated distribution (such as generating oval-shaped handles regardless of input or producing broken knobs), our method achieves consistent high-fidelity results that meet the physical requirements.
Table 1: Quantitative generation results on self-moving parts. “Phy.” denotes our physical plausibility metric.
Method Slider drawer Hinge door
CD $\downarrow$ F-score $\uparrow$ Phy. (%) CD $\downarrow$ F-score $\uparrow$ Phy. (%)
3D-CVAE 0.00248 0.683 38.9 0.00106 0.867 45.6
Ours (w/o loss guidance) 0.00082 0.818 49.3 0.00083 0.873 71.1
Ours (w/ loss guidance) 0.00060 0.862 74.3 0.00081 0.883 92.2
Table 2: Quantitative generation results on dependent parts. “Phy.” denotes our physical plausibility metric.
Method Line handle Hinge knob
CD $\downarrow$ F-score $\uparrow$ Phy. (%) CD $\downarrow$ F-score $\uparrow$ Phy. (%)
3D-CVAE 0.01000 0.403 63.3 0.00028 0.894 31.5
Ours (w/o loss guidance) 0.00174 0.742 71.1 0.00026 0.912 78.9
Ours (w/ loss guidance) 0.00173 0.744 85.9 0.00024 0.917 84.2
4.2 Downstream Applications
3D Printing. To better illustrate the physical plausibility of our generated parts, we 3D-printed the generated meshes. The printed parts not only fit seamlessly into the corresponding articulated objects but also demonstrated highly effective motions. Figure 6(a) shows these 3D-printed shapes, which were scaled down to $\frac{1}{200}$ of their original size, resulting in dimensions of approximately $10\text{cm}\times 10\text{cm}\times 10\text{cm}$.
Simulation for Robot Manipulation. Finally, we demonstrate that the generated parts can be manipulated using simple heuristics to validate the physical plausibility and quality of our generated parts. Following the setup in [17, 16], we load the objects in URDF format into the IsaacGym simulator [41] and apply a heuristic policy based on the estimated pose. Fig. 6(c) shows that our generated shapes exhibit high quality and can be successfully manipulated.
Table 3: Ablation studies on the pose proposal model and the conditional generation model on self-moving parts.
Method Slider drawer Hinge door
CD $\downarrow$ F-score $\uparrow$ Phy. (%) CD $\downarrow$ F-score $\uparrow$ Phy. (%)
Regress. pose proposal 0.00059 0.847 73.6 0.00134 0.736 50.0
w/o bbox condition 0.00178 0.740 53.5 0.00433 0.791 74.4
Full model 0.00060 0.862 74.3 0.00081 0.883 92.2
4.3 Ablation Studies and Analysis
This section delves into the effectiveness of different model components through ablation studies. Please refer to Appendix 0.B.1 for additional results and analysis,
Bounding Box Condition. We first verify the design of adding the bounding box condition to the diffusion model. As shown by the results in Tabs. 3 and 4, and Fig. 9, relying solely on the point cloud condition often yields part generation results that fail to fit the input object, such as drawers colliding or doors not filling the whole space. Conversely, incorporating the bounding box as an auxiliary condition significantly improves the overall performance of our model.
Pose Proposal Model. We then verify the effectiveness of adopting the diffusion-based pose proposal model. As shown in Tab. 3, leveraging the diffusion-based structure produces object bounding boxes of higher accuracy, thus benefiting the overall part generation quality.
Discussion on Generation Diversity. Using both the object point cloud and predicted part bounding box inherently restricts the diversity of parts that can be generated. This phenomenon is particularly notable for self-moving part categories, which must precisely conform to the input object’s specifications in size and shape. However, for dependent parts like handles and knobs, our model demonstrates the capacity to generate multiple physically plausible part instances, as illustrated in Fig. 7.
Figure 7: Illustration of part generation diversity. Despite the strict requirements imposed by the input conditions, our method could generate multiple plausible instances for dependent parts.
5 Conclusions
In this paper, we address the critical challenge of generating physically plausible parts for interactable objects. We propose a diffusion-based framework that integrates geometric conditioning and physical constraints, ensuring high accuracy and fidelity of the generated parts in shape and function under real-world interactions. Our results demonstrate substantial improvements in shape quality and physical plausibility over existing methods. These enhancements significantly benefit downstream applications such as 3D printing and robot manipulation, where precision is crucial for the tasks’ success.
Limitations and Future Work. A major limitation is that we only studied articulated objects from common datasets with 1D revolute or prismatic joints. Further studies can be done on more complex joints such as screws and wrenches with joint motions represented by Plücker coordinates. Such complex joint types can also be coupled with complex part geometries, which are important to certain domains such as tools or mechanical structures. We believe future studies in this direction will facilitate a variety of downstream applications such as CAD design and mechanical engineering.
Appendix 0.A Implementation Details
0.A.1 Physical-Aware Losses and Metric
Physical-Aware Losses. We scale all parts to the unit cube with padding $0.2$ before transforming into SDF. The threshold $\alpha$ for both physical-aware losses is $0.005$.
Physical-Plausible Metric. We detail the methodology of our proposed physical plausibility metric as follows. We first preprocess the meshes and discard those not watertight or have multiple large connected parts deemed physically infeasible. For the viable meshes, we employ a grid search with a grid length of $0.005$ around the part’s ground truth position to identify all positions where both collision loss (Eq. 5) and contact loss (Eq. 6) fall below a predefined threshold of $\beta=0.04$. A part is then considered physically plausible if moving it along the $-z$ axis for a distance of $0.02$ (applicable to hinge doors) or along the $+x$ and $-x$ axes for a distance of $0.04$ (other parts categories) results in the losses exceeding the threshold. Note that the hyperparameters are chosen to ensure that more than 95% of each category’s ground truth meshes are physically feasible. The pseudo-code for this algorithm is provided in Alg. 1.

### v3

# PhysPart: Physically Plausible Part Completion for Interactable Objects

@@ Abstract
I Introduction
The creation of 3D objects that facilitate human interactions is a critical and long-standing problem in industrial engineering, with applications ranging from manufacturing to the development of simulation environments. With the recent proliferation of AI-facilitated 3D content creation [48, 32, 10, 65, 78, 37, 51, 36], a line of efforts has been made to model interactable objects, particularly articulated ones, with 3D generative models [28, 35]. Imagine having a cabinet with a broken drawer; rather than engaging in labor-intensive carpentry with jigs and saws, a more efficient solution would be to generate a 3D model of the drawer and print it using a 3D printer, thus eliminating the need for specialized human expertise.
However, unlike rigid objects, whose modeling primarily focuses on shapes and appearances, interactable object modeling presents the additional challenge of intricate part motions under human interactions, constrained by geometry and physics. This complexity is particularly problematic for real-world applications, where high accuracy and robustness are essential. For instance, in the case of a drawer, even a slight size mismatch or a minor surface bump can cause it to get stuck in the cabinet.
In this paper, we address the problem of generating 3D object parts that are physically plausible both in their rest states and under interactions. Previous work on physics-constrained shape generation has predominantly concentrated on rigid objects, emphasizing connectivity [40, 21, 22] or structural stability [39]. However, for interactable objects with part motions, a significant challenge arises: while the generated shapes are static, the inherent state changes are crucial for physical plausibility. Consequently, both “first-order” constraints on static states and “second-order” constraints on state changes must be carefully considered.
For evaluation, we construct a benchmark with metrics for assessing both the shapes and the physics of the generation results. Besides the widely adopted metrics for 3D generative models, we introduce a physical plausibility metric that directly evaluates the part motions. With both qualitative and quantitative results, we show that our framework generates high-quality shapes, outperforming methods that lack proper modeling of geometric conditioning or physical constraints. Furthermore, we demonstrate 3D printing and robot manipulation applications, showcasing our strong potential in real-world applications that require high physical plausibility.
To summarize, our key contributions are:
  • We introduce the task of physically plausible part completion for interactable objects, along with a benchmark that includes comprehensive evaluation metrics for both shape and physical properties.
  • We propose a diffusion-based part generative model that incorporates geometric conditioning and physical constraints, formulated as stability and mobility losses. Our model achieves superior results in both shape quality and physical plausibility compared to baseline models, demonstrating its effectiveness.
  • We highlight physics-intensive applications of our model, such as 3D printing, robotic manipulation, and sequential part generation, where high physical realism is essential.
II Related Work
Articulated Object Modeling. Articulated object modeling is a crucial and longstanding field in 3D vision and robotics, encompassing a wide range of work in perception [16, 68, 11, 30, 33, 15, 17, 14, 18], reconstruction [2, 43, 24, 62], and generation [28, 35]. These contributions have greatly advanced applications in simulation [18, 67] and robot manipulation [66, 4, 12, 69, 29, 14, 15]. Beyond these perspectives, the intricate per-part state changes and inter-part motions, which formulates how objects interact and respond to interactions, also play important roles. For instance, [16] categorizes parts based on their actions, [68] learns part segmentation by analyzing point clouds during part motions, and [43, 24, 62] focus on reconstructing parts and joints across various motion states. Most studies focus on part motions at the pose level, often overlooking the crucial interaction between physics and detailed part geometry. This gap is particularly important in 3D generation, where physical plausibility is key for practical use. However, recent research has under-explored this aspect, concentrating more on structure and shape [28, 35]. For a more comprehensive overview, refer to [34].
Physics-Aware Shape Generation. Previous research in physics-aware 3D object generation has primarily focused on rigid shapes without complex self-interactions. [40, 21, 22] have used persistence diagrams to ensure topological connectivity, and [39] introduced a differentiable physical simulation layer to enhance stability under gravity. In contrast, physical principles have been extensively applied in other areas, such as in PhyScene [67], which aids in synthesizing interactable 3D scenes, and in human pose and motion generation, where studies [74, 75, 63] incorporate physics-aware constraints to reduce collisions and enhance realism. Additionally, built on diffusion models, SceneDiffuser [23] and PhysDiff [71] achieve physically plausible generation by incorporating physical optimization into the denoising process. Our approach extends this by embedding physical constraints into a generative diffusion model, enabling the creation of physically plausible parts.
3D Shape Completion. Shape completion is a vital step for reconstructing the missing parts of 3D shapes, traditionally tackled with techniques such as Laplacian hole filling [57, 45, 76] and Poisson surface reconstruction [25, 26]. These methods primarily address small gaps and basic geometric forms. Another strategy employs structural regularities, like symmetries, to infer unobserved parts of shapes [61, 41, 47, 52, 58]. With the advent of extensive 3D datasets, retrieval-based approaches [59, 31, 44, 27] have emerged, searching databases to find the best matches for incomplete inputs, alongside learning-based methods [46, 13, 9, 7, 70, 19, 56, 6] that minimize discrepancies between network predictions and actual shapes. Notably, 3D-EPN [9] utilizes a 3D encoder-decoder architecture to predict complete shapes from partial volumetric data, while Scan2Mesh [8] transforms range scans into 3D meshes through direct mesh surface optimization. PatchComplete [50] leverages local structural priors to fill in shapes from unseen categories. Alternatively, generative methods like GANs [77, 73, 3, 64, 53] and AutoEncoders [42, 1] offer a different approach by generating diverse, plausible shapes from partial inputs, though they often compromise on completion accuracy. Our method leverages diffusion models and physics-aware loss guidance, effectively reducing surface artifacts and surpassing current state-of-the-art methods in producing realistic and physics-plausible 3D shapes.
III Method
Given the full point cloud of an articulated object $\mathcal{O}$ that has a missing part, we aim to generate a replacement part $p$ that integrates seamlessly into the object while adhering to geometric and physical constraints. We employ volumetric Truncated Signed Distance Fields (T-SDF) to model object parts and train a 3D-VQVAE to compress these models into a discrete and compact latent space. Unlike prior research that employs SDF representation that often overlooks the object’s scale as generating SDF usually involves normalizing the part to the unit cube, our framework also emphasizes accurate scaling of the objects for manufacturing purposes. Upon this part representation, our method utilizes a coarse-to-fine approach, involving two diffusion-based sub-models: a pose proposal model that determines the part’s approximate position at the coarse level, and a part code generator that refines the shape based on geometric conditions (Sec. III-A) and physical constraints (Sec. III-B). This integrated framework ensures that the generated part can interact properly with external forces and fit accurately within the existing object structure. An overview of our pipeline is shown in Fig. 1.
Part-Pose Proposal. We propose a diffusion-based pose estimator to determine the bounding box of the part to be generated from the object point cloud $\mathcal{P}$. Drawing inspiration from GenPose [72], we train a score-based diffusion model $\Phi_{\theta}$ alongside an energy-based diffusion model $\Psi_{\phi}$. First, we generate pose candidates ${p^{i}}{i=1}^{K}$ using $\Phi\theta$. Then, the pose energies $\Psi_{\phi}(p^{i},\mathcal{P})$ are computed for each candidate via $\Psi_{\phi}$. By ranking the candidates based on their energies and filtering out the lower-ranked ones, the remaining bounding box candidates are aggregated into the final output using mean pooling.
TABLE I: Quantitative generation results on self-moving parts. “Phys.” denotes our percentage physical plausibility metric. “Ours, w/o loss-guided samp.” denotes our method without using the proposed loss-guided sampling.
Self-moving parts Dependent parts
Method Slider drawer Hinge door Line handle Hinge knob
@@ III Method
C-VAE 0.00248 0.683 38.9 0.00106 0.867 45.6 0.01000 0.403 63.3 0.00028 0.894 31.5
Ours, w/o loss-guided samp. 0.00082 0.818 49.3 0.00083 0.873 71.1 0.00174 0.742 71.1 0.00026 0.912 78.9
Ours 0.00060 0.862 74.3 0.00081 0.883 92.2 0.00173 0.744 85.9 0.00024 0.917 84.2
III-A Geometry-Conditioned Part Generation
SDF Generation with Latent Diffusion.
@@ III-A Geometry-Conditioned Part Generation
SDFusion [5] introduces a diffusion-based 3D generative model. The approach begins by compressing high-dimensional SDFs into a compact latent space using a 3D-VQVAE. Specifically, an SDF $X\in\mathbb{R}^{D\times D\times D}$ is encoded into a latent vector $z\in\mathbb{R}^{d\times d\times d}$ by the encoder $E_{\phi}$. The encoded SDF is then reconstructed by the decoder $D_{\tau}$ after passing $z$ through a vector quantization codebook VQ, i.e.,
@@
Afterwards, conditional diffusion model is trained on this learned latent distribution. The classifier-free guidance [20] and DDIM [55] sampling techniques are employed to enhance the flexibility and efficiency of the generation process.
Geometric Conditioning. The above framework allows the generation of missing parts using point cloud as the condition. However, due to limited training data and the point cloud’s inherent ambiguity, relying solely on the point cloud often yields unsatisfactory results. To improve the model’s performance, we leverage the part bounding box predicted by our pose estimation model. Specifically, given the predicted bounding box of the desired part, we rotate to canonical space and employ the ratio of the dimensions along the three axes as an auxiliary condition for the diffusion model. We use the length ratio rather than absolute values since SDF representation is scale-agnostic. Note that although we allow rotation to the input point cloud, the mesh is always generated in its canonical pose.
Empirically, we adopt the PointNet++ [49] ($\phi_{1}$) to encode the point cloud $c_{1}\in\mathbb{R}^{N\times 3}$, and a sinusoidal positional encoding followed by a two-layer MLP ($\phi_{2}$) to encode the bounding box ratio $c_{2}\in\mathbb{R}^{3}$. The training objective for the diffusion model $\theta$ and encoders $\phi_{1},\phi_{2}$ is formulated as:
$L(\theta,$ $\phi_{1},\phi_{2}):=$
$\mathbb{E}_{z,c,e,t}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t},t,D(\phi_{1}(c_{1})),D(\phi_{2}(c_{2})))\right\|^{2}\right],$ (2)
where $D(\cdot)$ denotes a dropout operation, $c_{1},c_{2}$ are the conditions, and $z_{t}$ is the noisy latent variable at timestamp $t$.
@@
III-B Loss-Guided Physical Constraints
Given an object’s point cloud, we predict the bounding box for the missing part, generate the desired part by sampling from the diffusion model using both the object’s point cloud and the bounding box as conditions, and scale the generated mesh using the bounding box volume ratio. However, this generation process is agnostic to physical constraints and may result in parts incompatible with the original object. To address this, we introduce physics-aware losses to guide the diffusion sampling process, ensuring physical compatibility.
Stability Losses. The most critical criterion for the generated part is to avoid collision with the provided object point cloud $\mathcal{P}$. Therefore, we define the collision loss by considering the predicted part-SDF value $X(p)$ of the object point $p$:
$\ell_{coll}(X,\mathcal{P})=\sum_{p\in\mathcal{P}}\text{ReLU}(-X(p)-\alpha),$ (3)
@@ III-B Loss-Guided Physical Constraints
$\ell_{cont}(X,\mathcal{P})=\min_{p\in\mathcal{P}}\text{ReLU}(X(p)-\alpha),$ (4)
@@
TABLE II: Ablation studies on the pose proposal model and conditional generation model.
Self-moving parts Dependent parts
Method Slider drawer Hinge door Line handle Hinge knob
[…]
Regress. pose proposal 0.00059 0.847 73.6 0.00134 0.736 50.0 0.00252 0.655 74.2 0.00036 0.847 81.1
w/o bbox condition 0.00178 0.740 53.5 0.00433 0.791 74.4 0.00784 0.572 67.0 0.00033 0.885 80.7
Full model 0.00060 0.862 74.3 0.00081 0.883 92.2 0.00173 0.744 85.9 0.00024 0.917 84.2
@@
$\ell_{coll-m}(X,\mathcal{P})=\mathbb{E}_{\hat{X}\sim M(X)}\sum_{p\in\mathcal{P}}\text{ReLU}(-\hat{X}(p)-\alpha),$ (5)
$\ell_{cont-m}(X,\mathcal{P})=\mathbb{E}_{\hat{X}\sim M(X)}\min_{p\in\mathcal{P}}\text{ReLU}(\hat{X}(p)-\alpha).$ (6)
@@
Loss-Guided Sampling. Conditional diffusion models aim to generate samples $x_{0}$ given the condition $y$, where the conditional score at timestamp $t$ is derived using Bayes’ rule:
@@
$\overline{\epsilon}_{\theta}(x_{t},t,y)=\epsilon_{\theta}(x_{t},t)-\sqrt{1-\overline{\alpha}_{t}}\omega\cdot\nabla_{x_{t}}\log{p_{t}}(y|x_{t}),$ (8)
where the gradient term $\nabla_{x_{t}}\log{p_{t}}(y|x_{t})$ is written as $-\frac{1}{\sqrt{1-\overline{\alpha}_{t}}}(\epsilon_{\theta}(x_{t},t,y)-\epsilon_{\theta}(x_{t},t))$ in classifier-free guidance.
To introduce physical constraints into the sampling process, we approximate the gradient using our defined physics-aware loss functions. As loss functions are defined on noise-free samples, we compute the loss on the predicted noise-free sample at time step $t$, denoted as $\hat{x}_{0}$:
$\nabla_{x_{t}}\log{p_{t}}(y|x_{t})$ $\approx\nabla_{x_{t}}\log{p_{t}}(y|\hat{x}_{0})$
$=\nabla_{x_{t}}\log\frac{\exp(-\ell_{y}(\hat{x}_{0}))}{Z}=-\ell_{y}(\hat{x}_{0}),$ (9)
where $Z$ is the normalizing constant independent of $x_{t}$.
Thus, we combine classifier-free guidance and loss guidance to obtain the modified score:
@@
$\ell_{y}(x_{0})=\ell_{coll}+\ell_{cont}+\ell_{coll-m}+\ell_{cont-m}.$ (11)